# Volatility Parametrizations with Random Coefficients: Analytic Flexibility for Implied Volatility Surfaces

### v2

@@ 2.2 Volatility Parametrizations with Random Parameters
The map $\hat{\sigma}(T,K)$ as defined Equation 2.8 is thus an arbitrage-free implied volatility surface, and the pricing of European-style options will collapse to determining a mixture distribution weighted with the probability distribution of $\vartheta$. If we specify the probability distribution of the random variable $\vartheta$ in parametric form, we can combine the sets of the parametrized surface and the random variable to a common set of parameters. Suppose that the random variable $\vartheta$ is given in parametric form by the parameters $\overline{q}=(q_{1},q_{2},\dots,q_{l})\in\mathbb{R}^{l}$. We combine the parameter sets of $\overline{q}$ and $\overline{p}$ to an extended parameter parameter vector
@@
As the pricing surface of the randomized parametrization $V_{c/p}(T,K;\overline{p}^{*})$ is defined as an integral over the domain of parameters, weighted by the PDF of the random variable $\vartheta$, we discretize the integral using a numerical integration scheme. We have to ensure that the discretization is arbitrage-free, for which we utilize an integration technique that maintains the convex property of the summation. The Gaussian quadrature integration approximates an integral of the form $\int_{a}^{b}h(x)w(x){\rm d}x$, on a real domain $[a,b]\subset\mathbb{R}$, for an integrable function $h(x)$ and a weight function $w(x)$, as the sum
@@
where $\{\lambda_{n},\theta_{n}\}_{n=1}^{N_{q}}$ are called Gauss-quadrature weights. The important property of the Gauss-quadrature integration is that the integral is exact for any polynomial of degree less than $2N_{q}$. Since the expression $h(x)=1$ is a polynomial of degree 00, the approximation is exact for this integrate and we obtain
$\int_{a}^{b}1w(x){\rm d}x=\int_{a}^{b}w(x){\rm d}x=\sum_{n=1}^{N_{q}}\lambda_{n}.$ (2.11)
The weights of a Gaussian quadrature approximation thus always sum 1. If we apply the integration technique to the randomized price surface with $w(x)=f_{\vartheta}(x)$, we obtain that
$V_{c/p}(T,K;\overline{p}^{*})=\int_{\mathbb{R}}V_{c/p}(T,K;\overline{p}(\theta))f_{\vartheta}(\theta){\rm d}\theta\approx\sum_{n=1}^{N_{q}}\lambda_{n}V_{c/p}(T,K;\overline{p}(\theta_{n})),$ (2.12)
where the pairs $\{\lambda_{n},\theta_{n}\}_{n=1}^{N_{q}}$ are the Gauss-quadrature weights and nodes. These weights and nodes depend on the distribution function $F_{\vartheta}(\cdot)$ of $\vartheta$ and can be calculated explicitly as outlined in A. We define the truncated, or discretized pricing surface with $N_{q}$ terms as the sum
@@
Since the discretization is only an approximation of $V_{c/p}(T,K;\overline{p}^{*},N_{q})$ we have to ensure that it is arbitrage-free too.
@@
Let $\{\lambda_{n},\theta_{n}\}_{n=1}^{N_{q}}$ be the Gauss-quadrature pairs of $\vartheta$ of degree $N_{q}$ and define the function $f_{\bar{\vartheta}}(\cdot)\colon\mathbb{R}\to\mathbb{R}_{+}$ such that
@@
We claim that this is a probability mass function. Since $\lambda_{n}\geq 0$ for all $n\leq N_{q}$, the function is certainly positive. Furthermore, since quadrature integration is exact for polynomials up to $2N_{q}$ degree, we have as in (2.11), that $\int_{\mathbb{R}}f_{\bar{\vartheta}}(x){\rm d}x=1$. By Skorohod’s representation theorem, there exists a random variable $\bar{\vartheta}$, such that $f_{\bar{\vartheta}}$ is its probability mass function. It follows that (2.14) is true, since $\mathbb{E}[V_{c/p}(T,K;\overline{p}(\bar{\vartheta}))]$ is given by $\sum_{n=1}^{N_{q}}\lambda_{n}V_{c/p}(T,K;\overline{p}(\theta_{n}))$. ∎
Remark (Choice of $N_{q}$).
Generally, a larger choice of $N_{q}$ means that the discretized randomization approximates the continuous randomization better. However, since the discretization is arbitrage-free for any $N_{q}\in\mathbb{N}$, an exact approximation of a continuous $\vartheta$ is not necessary, and $N_{q}$ can be treated simply as a parameter to further specify the nature of the randomization of the parameters.
@@ 2.3 Analytic Expansion of Randomized Volatility Surface
With the additional parameter and the help of the quadrature integration, we transformed the semi-analytical pricing surface into a finite sum of prices given by the discretized randomization of (2.13). Since the Black-Scholes equation has no analytical inverse in the volatility argument, the implied volatility surface $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$ cannot be obtained explicitly. Nevertheless, Brigo and Mercurio  [4] showed how to get a polynomial expansion of the implied volatility function based on the Taylor expansion under a mixture model of lognormal prices. We derive a generalization of this result, and we use it to obtain an analytic expression of $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$. Let $(T,K)\in\Pi$ be fixed and suppose we want to find the value of $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$. We first define the function $m(T,K)$ as
@@
$f(m,P(m)):=S_{0}\bigg{[}\Phi\bigg{(}\frac{m+\frac{1}{2}\hat{\sigma}^{2}(m)T}{P(m)\sqrt{T}}\bigg{)}-{\rm e}^{-m}\Phi\bigg{(}\frac{m-\frac{1}{2}\hat{\sigma}^{2}(m)T}{P(m)\sqrt{T}}\bigg{)}\bigg{]}.$ (2.16)
@@
If we equate the two equations, since $P(m(T,K))$ is equal to $\hat{\sigma}(T,K,\overline{p}(\theta_{n}))$, the equation
$f(m,\hat{\sigma}(m))=g(m),$ (2.18)
is exactly equal to (2.13) at $m=m(T,K)$. The goal is to obtain a polynomial expansion of the function $P(m)$, which we can then evaluate at $m=m(T,K)$ to obtain the randomized volatility function at $(T,K)$. The expansion utilizes the ideas from the implicit function theorem [17] to obtain any higher-order derivative from an implicit function, such as Equation 2.18. We can obtain the derivatives of $P(m)$ by differentiating both sides with respect to $m$ and write the function $P(m)$ as a Taylor expansion in terms of the derivatives at 00 and the function $P(m)$ at $m=0$.
@@
be the log-moneyness with interest-rate $r$. Then, the discretized randomization $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$ at $(T,K)$ is given by the Taylor expansion $P_{(T,K)}(m)\colon\mathbb{R}\to\mathbb{R}$, evaluated at $m=m(T,K)$, where
@@
$\qquad\quad+\sum_{n=1}^{N_{q}}\lambda_{n}\bigg{[}\frac{E_{n}}{H_{n}^{5}}\big{(}3-16H_{n}^{2}+31H_{n}^{4}\big{)}\bigg{]}\Bigg{\}},$
@@
$\Sigma_{0}:=\frac{1}{2}P_{(T,K)}(0)\sqrt{T},\qquad\Sigma_{2}:=P_{(T,K)}(0)P_{(T,K)}^{(2)}(0)T,\qquad\Sigma_{4}:=P_{(T,K)}(0)P_{(T,K)}^{(4)}(0)T,$
$H_{n}:=\frac{1}{2}\eta_{n}\sqrt{T},\qquad E_{n}:=\exp{\Big{(}\frac{1}{2}(\Sigma^{2}_{0}-H^{2}_{n})\Big{)}},\qquad\eta_{n}=\hat{\sigma}(T,K;\overline{p}(\theta_{n})).$
The expression $\mathcal{O}(m^{8})$ is the usual Big-O notation as $m\to 0$, meaning that for the remainder $R(m)$ we have $\lim_{x\to 0}\frac{R(m)}{m^{8}}\leq C$ for some constant $C$.
@@
We can now obtain any order derivative of $P_{(T,K)}(m)$ by differentiating both sides by $m$ and express the $n$-order derivatives of $P_{(T,K)}(m)$ as partial derivatives of $f$ of at most $n$ degrees. Note, however, that the number of derivatives on the right-hand side grows exponentially [^5] More precisely, the growth is exponentially bounded from above (by $2^{i-1}$ with $i>0$ the order of differentiation). It is possible to reorder and collect the terms in such a way the overall number is slightly lower, but in any case, the growth is “more than” polynomial. in the order of the target derivative (due to the formula of the derivation of a product). As it turns out, all the second-order derivatives disappear, simplifying the calculation significantly.
@@
$f\left(0,P_{(T,K)}(0)\right)=S_{0}\bigg{[}\Phi\bigg{(}P_{(T,K)}(0)\sqrt{T}\bigg{)}-\Phi\bigg{(}-P_{(T,K)}(0)\sqrt{T}\bigg{)}\bigg{]},$
@@
$g(0)=S_{0}\sum_{n=1}^{N_{q}}\lambda_{n}\bigg{[}\Phi\bigg{(}\eta_{n}\sqrt{T}\bigg{)}-\Phi\bigg{(}-\eta_{n}\sqrt{T}\bigg{)}\bigg{]},$
@@
Combining the results, we can express the function $P_{(T,K)}(m)$ as its Taylor expansion function around 00, which yields (2.19), and obtain the value for $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$ by evaluating at $m(T,K)$. ∎
Figure 1: The figure shows the randomization of a parametrization with $N_{q}=2$ for a fixed time $T$. The grey plots show the individual slices which are mixed, and the green lines show the function $P_{(T,K)}$ for three different $K$s. For each $P_{(T,K)}$, the randomized volatility meets the expansion $P_{(T,K_{i})}$ exactly at $K_{i}$.
@@ 3 Illustrative Examples
The process of randomization transforms a parametric implied volatility surface into a new parametric implied volatility surface with extra parameters and flexibility. We examine a few examples of this randomization process to better understand its mechanics and effectiveness.

### v3

@@ 2.2 Volatility Parametrizations with Random Parameters
The map $\hat{\sigma}(T,K)$ as defined Equation 2.9 is thus an arbitrage-free implied volatility surface, and the pricing of European-style options will collapse to determining a mixture distribution weighted with the probability distribution of $\vartheta$. If we specify the probability distribution of the random variable $\vartheta$ in parametric form, we can combine the sets of the parametrized surface and the random variable into a common set of parameters. Suppose that the random variable $\vartheta$ is given in parametric form by the parameters $\overline{q}=(q_{1},q_{2},\dots,q_{l})\in\mathbb{R}^{l}$. We combine the parameter sets of $\overline{q}$ and $\overline{p}$ to an extended parameter parameter vector
@@
As the pricing surface of the randomized parametrization $V_{c/p}(T,K;\overline{p}^{*})$ is defined as an integral over the domain of parameters, weighted by the law of the random variable $\vartheta$, we now derive a discretization of the Lebesgue integral using a numerical integration scheme. While the discretization acts an approximation of the distribution of $\vartheta$ and the pricing surface, we will ensure that the discretization remains arbitrage-free for any degree of accuracy of the approximation. This is crucial as the approximation will be the primary pricing surface parametrization to fit to the market, and the degree of approximation will act as an additional parameter to the surface. To ensure this property we choose a numerical integration scheme that maintains the convex property of the summation. The Gaussian quadrature is a suitable tool for such an approximation.
Suppose we wish to integrate an expression $\int_{a}^{b}h(x)w(x){\rm d}x$, on a real domain $a,b]\subset\mathbb{R}$, for an integrable function $h(x)$ and a weight function $w(x)$ (a positive function which integrates to 1). The Gaussian quadrature approximates this as
@@
where $\{\lambda_{n},\theta_{n}\}_{n=1}^{N_{q}}$ are known as the Gauss-quadrature weights. The important property of the Gauss-quadrature integration is that the integral is exact for any polynomial of degree less than $2N_{q}$. Since the expression $h(x)=1$ is a polynomial of degree 0, the approximation is exact for this integral for any $N_{q}$ [^5] Other integration methods do not have this property. For instance, a trapezoidal method with only two grid points will yield 0 if $w(x)$ has finite support., and we obtain
$\int_{a}^{b}1w(x){\rm d}x=\int_{a}^{b}w(x){\rm d}x=\sum_{n=1}^{N_{q}}\lambda_{n}=1.$ (2.12)
In particular, this means that for an absolutely continuous real-valued random variable $X$, where the probability density function $w(x)=f_{X}(x)$ is a weight function, we can approximate its expectation as
$\mathbb{E}[g(X)]=\int_{\mathbb{R}}g(x)f_{X}(x){\rm d}x\approx\sum_{n=1}^{N_{q}}\lambda_{n}g(x_{n})$ (2.13)
where the pairs $\{\lambda_{n},x_{n}\}_{n=1}^{N_{q}}$ are the Gauss-quadrature weights and nodes. The Gauss quadrature works more generally, as outlined in A, to approximate the expectation $\mathbb{E}[g(X)]$ for any real-valued random variable. Applied to the randomized surface, we find
$V_{c/p}(T,K;\overline{p}^{*})=\int_{\mathbb{R}}V_{c/p}(T,K;\overline{p}(\theta)){\rm d}\mu(\theta)\approx\sum_{n=1}^{N_{q}}\lambda_{n}V_{c/p}(T,K;\overline{p}(\theta_{n})),$ (2.14)
where the pairs $\{\lambda_{n},\theta_{n}\}_{n=1}^{N_{q}}$ are the Gauss-quadrature weights and nodes depend on the distribution function $F_{\vartheta}(\cdot)$ of $\vartheta$ (see A). We define the truncated, or discretized pricing surface with $N_{q}$ terms as the sum
@@
Since the discretization approximates $V_{c/p}(T,K;\overline{p}^{*},N_{q})$, and due to the convex property, it is arbitrage-free for any $N_{q}$.
@@
Let $\{\lambda_{n},\theta_{n}\}_{n=1}^{N_{q}}$ be the Gauss-quadrature pairs of $\vartheta$ of degree $N_{q}$ and define the function
@@
We claim that this is a probability mass function. Since $\lambda_{n}\geq 0$ for all $n\leq N_{q}$, the function is certainly positive. Furthermore, since quadrature integration is exact for polynomials up to $2N_{q}$ degree, we have as in (2.12), that $\sum_{n=1}^{N_{q}}\lambda_{n}=1$. By Skorohod’s representation theorem, there exists a random variable $\bar{\vartheta}$, such that $f_{\bar{\vartheta}}$ is its probability mass function. It follows that (2.16) is true, since $\mathbb{E}[V_{c/p}(T,K;\overline{p}(\bar{\vartheta}))]$ is given by $\sum_{n=1}^{N_{q}}\lambda_{n}V_{c/p}(T,K;\overline{p}(\theta_{n}))$. ∎
@@ 2.3 Analytic Expansion of Randomized Volatility Surface
With the additional parameter and the help of the quadrature integration, we transformed the semi-analytical pricing surface into a finite sum of prices given by the discretized randomization of (2.15). Since the Black-Scholes equation has no analytical inverse in the volatility argument, the implied volatility surface $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$ cannot be obtained explicitly. Nevertheless, Brigo and Mercurio  [brigo2002lognormal] showed how to get a polynomial expansion of the implied volatility function based on the Taylor expansion under a mixture model of lognormal prices. We derive a generalization of this result, and we use it to obtain an analytic expression of $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$. Let $(T,K)\in\Pi$ be fixed and suppose we want to find the value of $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$. We first define the function $m(T,K)$ as
@@
$f(m,P(m)):=S_{0}\bigg[\Phi\bigg(\frac{m+\frac{1}{2}P^{2}(m)T}{P(m)\sqrt{T}}\bigg)-{\rm e}^{-m}\Phi\bigg(\frac{m-\frac{1}{2}P^{2}(m)T}{P(m)\sqrt{T}}\bigg)\bigg].$ (2.18)
@@
If we equate the two equations, the equation
$f(m,P(m))=g(m),$ (2.20)
is exactly equal to (2.15) at $m=m(T,K)$. The goal is to obtain a polynomial expansion of the function $P(m)$, which we can then evaluate at $m=m(T,K)$ to obtain the randomized volatility function at $(T,K)$. The expansion utilizes the ideas from the implicit function theorem [krantz2002implicit] to obtain any higher-order derivative from an implicit function, such as Equation 2.20. We can obtain the derivatives of $P(m)$ by differentiating both sides with respect to $m$ and write the function $P(m)$ as a Taylor expansion in terms of the derivatives at 0 and the function $P(m)$ at $m=0$. Prior to stating the full expansion, we show a property of the implicit equation, which simplifies the expansion terms.
Lemma 2.4.
Let the functions $f$ and $g$ be defined as Equation 2.18 and Equation 2.19. The set $S=\{(x,y)\in\mathbb{R}^{2}:f(x,y)=g(x)\}$ is symmetric over the $y$-axis.
Proof.
We will show directly that if $(x,y)\in S$, then the point (x,y)(-x,y) is also contained in $S$. Assuming $(x,y)\in S$, we have
$\Phi\bigg(\frac{x+\frac{1}{2}yT}{y\sqrt{T}}\bigg)-{\rm e}^{-x}\Phi\bigg(\frac{x-\frac{1}{2}yT}{y\sqrt{T}}\bigg)=\sum_{n=1}^{N_{q}}\lambda_{n}\bigg[\Phi\bigg(\frac{x+\frac{1}{2}\eta_{n}^{2}T}{\eta_{n}\sqrt{T}}\bigg)-{\rm e}^{-x}\Phi\bigg(\frac{x-\frac{1}{2}\eta_{n}^{2}T}{\eta_{n}\sqrt{T}}\bigg)\bigg].$
Multiplying both sides with $e^{x}$ and adding a term $1-e^{x}$ on both sides we find
$1-{\rm e}^{x}+{\rm e}^{x}\Phi\bigg(\frac{x+\frac{1}{2}yT}{y\sqrt{T}}\bigg)-\Phi\bigg(\frac{x-\frac{1}{2}yT}{y\sqrt{T}}\bigg)=1-{\rm e}^{x}+\sum_{n=1}^{N_{q}}\lambda_{n}\bigg[{\rm e}^{x}\Phi\bigg(\frac{x+\frac{1}{2}\eta_{n}^{2}T}{\eta_{n}\sqrt{T}}\bigg)-\Phi\bigg(\frac{x-\frac{1}{2}\eta_{n}^{2}T}{\eta_{n}\sqrt{T}}\bigg)\bigg].$
On the other hand, for (x,y)(-x,y), we have
$\frac{1}{S_{0}}f(-x,y)$ $=\Phi\bigg(\frac{-x+\frac{1}{2}yT}{y\sqrt{T}}\bigg)-{\rm e}^{x}\Phi\bigg(\frac{-x-\frac{1}{2}yT}{y\sqrt{T}}\bigg)$
$=1-e^{x}-\Phi\bigg(\frac{x-\frac{1}{2}yT}{y\sqrt{T}}\bigg)+{\rm e}^{x}\Phi\bigg(\frac{x+\frac{1}{2}yT}{y\sqrt{T}}\bigg)$
as well as
$\frac{1}{S_{0}}g(-x)$ $=\sum_{n=1}^{N_{q}}\lambda_{n}\bigg[\Phi\bigg(\frac{-x+\frac{1}{2}\eta_{n}^{2}T}{\eta_{n}\sqrt{T}}\bigg)-{\rm e}^{x}\Phi\bigg(\frac{-x-\frac{1}{2}\eta_{n}^{2}T}{\eta_{n}\sqrt{T}}\bigg)\bigg]$
$=1-e^{x}+\sum_{n=1}^{N_{q}}\lambda_{n}\bigg[-\Phi\bigg(\frac{x-\frac{1}{2}\eta_{n}^{2}T}{\eta_{n}\sqrt{T}}\bigg)+{\rm e}^{x}\Phi\bigg(\frac{x+\frac{1}{2}\eta_{n}^{2}T}{\eta_{n}\sqrt{T}}\bigg)\bigg]$
Equating the two terms we obtain that
$\frac{1}{S_{0}}f(-x,y)=\frac{1}{S_{0}}g(-x)\iff f(-x,y)=g(-x),$
and we therefore conclude that $(-x,y)\in S$. ∎
From the lemma we make the useful conclusion that implicit function $P(m)$ is even, and has therefore vanishing odd expansion coefficients.
@@
be the log-moneyness with interest-rate $r$. Then, the discretized randomization $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$ at $(T,K)$ is given by the Taylor expansion $P_{(T,K)}(m)\colon A\subset\mathbb{R}\to\mathbb{R}$, evaluated at $m=m(T,K)$, where
@@
[…]
@@
$\Sigma_{0}=\frac{1}{2}P_{(T,K)}(0)\sqrt{T},\qquad\Sigma_{2}=P_{(T,K)}(0)P_{(T,K)}^{(2)}(0)T,\qquad\Sigma_{4}=P_{(T,K)}(0)P_{(T,K)}^{(4)}(0)T,$
$H_{n}=\frac{1}{2}\eta_{n}\sqrt{T},\qquad E_{n}=\exp{\Big(\frac{1}{2}(\Sigma^{2}_{0}-H^{2}_{n})\Big)},\qquad\eta_{n}=\hat{\sigma}(T,K;\overline{p}(\theta_{n})).$
The expression $\mathcal{O}(m^{8})$ is the usual Big-O notation as $m\to 0$ and the subset $A$ is the radius of convergence of the expansion. This means that for the remainder $R(m)$ we have $\lim_{m\to 0}\frac{R(m)}{m^{8}}\leq C$ for some constant $C$.
@@
We can now obtain any order derivative of $P_{(T,K)}(m)$ by differentiating both sides by $m$ and express the $n$-order derivatives of $P_{(T,K)}(m)$ as partial derivatives of $f$ of at most $n$ degrees. Note, however, that the number of derivatives on the right-hand side grows exponentially [^6] More precisely, the growth is exponentially bounded from above (by $2^{i-1}$ with $i>0$ the order of differentiation). It is possible to reorder and collect the terms in such a way the overall number is slightly lower, but in any case, the growth is “more than” polynomial. in the order of the target derivative (due to the formula of the derivation of a product). As we show in 2.4, the function $P_{(T,K)}$ is even, and therefore, all odd-order derivatives vanish, simplifying the calculation significantly.
@@
$f\left(0,P_{(T,K)}(0)\right)=S_{0}\bigg[\Phi\bigg(\frac{1}{2}P_{(T,K)}(0)\sqrt{T}\bigg)-\Phi\bigg(-\frac{1}{2}P_{(T,K)}(0)\sqrt{T}\bigg)\bigg],$
@@
$g(0)=S_{0}\sum_{n=1}^{N_{q}}\lambda_{n}\bigg[\Phi\bigg(\frac{1}{2}\eta_{n}\sqrt{T}\bigg)-\Phi\bigg(-\frac{1}{2}\eta_{n}\sqrt{T}\bigg)\bigg],$
@@
Combining the results, we can express the function $P_{(T,K)}(m)$ as its Taylor expansion function around 0, which yields (2.21), and obtain the value for $\hat{\sigma}(T,K;\overline{p}^{*},N_{q})$ by evaluating at $m(T,K)$. ∎
Figure 1: The figure shows the randomization of a parametrization with $N_{q}=2$ for a fixed time $T$. The grey plots show the individual slices that are mixed, and the green lines show the function $P_{(T,K)}$ for three different $K$s. For each $P_{(T,K)}$, the randomized volatility meets the expansion $P_{(T,K_{i})}$ exactly at $K_{i}$.
@@ 3 Illustrative Examples
The process of randomization transforms a parametric implied volatility surface into a new parametric implied volatility surface with extra parameters and flexibility. We examine a few examples of this randomization process to better understand its mechanics and effectiveness. A repository with a Python implementation is available on GitHub.